OpenStudy (anonymous):

how to plot -3y=-2x-6

OpenStudy (tkhunny):

Usually, a Slope-Intercept form can be used. I like the Intercept form: 2x - 3y = -6 ==> \(\dfrac{x}{-3}+\dfrac{y}{2} = 1\)

OpenStudy (mathmale):

Here's how to use this "intercept form:" First, let x=0. Solve for y. This gives you the point (0, y). Plot this point. Next, let y=0. Solve for x (and so on). Please do your drawing via the Draw utility (below). Draw a line thru the 2 points.

OpenStudy (tkhunny):

Nah, you just look at it and draw your line. No algebra. You can substitute and solve in ANY form. The Intercept Form is much more convenient. It also works nicely in 3D.

OpenStudy (akausen):

I think it's easiest in slope-intercept form, in which you solve for y. This would give you y=2/3x+2. b is the y-intercept, which is 2 since y=mx+b. Then use 2/3 to plot the rest, which is the slope, rise over run.

OpenStudy (tkhunny):

Well, sure. Lot's of folks think that's easiest because they are not familiar with alternatives.

OpenStudy (akausen):

I'm just trying to provide the easiest way because someone trying to learn doesn't need to start complicated.

OpenStudy (tkhunny):

I understand completely. The easiest way is always the way you know until you know an easier way. There is nothing to defend.